OpenStudy (anonymous):

Calculate the derivatives \(sin^{'}(x) = \frac{d}{dx}sin(x) and cos^{'}(x) = \frac{d}{dx}cos(x)\) by differentiating the summands of the power series. Show as a corollary the identity \(sin^{2}(x)+cos^{2}(x) = 1\) for all \(x \in \mathbb{R}\).

OpenStudy (anonymous):

\[\sin^{'}(x) = \frac{d}{dx}\sin(x) and \cos^{'}(x) = \frac{d}{dx}\cos(x)\]

OpenStudy (anonymous):

ah ok you want to write the power series, which is pretty easy, and then take the derivative of each of the terms then simplify and see what the power series now resembles

OpenStudy (anonymous):

Well it comes out to be cos x for the first one -sin x for the second one

OpenStudy (anonymous):

i it possible to solve it ? i am quite bad at mathematics.. :(

OpenStudy (anonymous):

@tunahan look at what i just said

OpenStudy (anonymous):

ok thx..

OpenStudy (anonymous):

if you need more help just ask, k?

OpenStudy (anonymous):

also it would help me more if i would see the solution, and try too understand it, because tomorrow i must give up this homework, and there is 3 more questions, i dont know if i can finish it until give up time... there is 5 more questions to do .

OpenStudy (anonymous):

well do you know how to write a power series?

OpenStudy (anonymous):

i am sorry as i said my math is really sucks, i am computer science student and i had always problem with math...

OpenStudy (anonymous):

http://www.khanacademy.org/math/calculus/v/sine-taylor-series-at-0--maclaurin watch this first then

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

http://www.khanacademy.org/math/calculus/v/taylor-polynomials or this and then write the taylor series

OpenStudy (anonymous):

@cerezas i have looked , but it will take me time understand it hole, i am in germany and in about 8 hours i must finish my homework :( its better if you can solve it for me, and in one month i have exam i will ask you every details of it if you will be here ;)

OpenStudy (anonymous):

i need 16 points if i have less then 8 points i think i cant participate at final exam

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=taylor+series+sin+x&lk=3 now take the derivative, simplify, and that's cosx

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

sure